OpenStudy (sierra!):

The length of a rectangular field is 4 m longer than its width. If the perimeter is 56 m, what is the width of the field? A. 8 m B. 12 m C. 14 m D. 16 m

OpenStudy (sierra!):

@Data_LG2 @iGreen

OpenStudy (sierra!):

@iGreen

OpenStudy (sierra!):

@sammixboo

OpenStudy (sierra!):

@sleepyjess

OpenStudy (sierra!):

@ksanka

OpenStudy (anonymous):

label the width a and so the length will be a+4

OpenStudy (anonymous):

then P=2(length+width)

OpenStudy (anonymous):

56=2(a+a+4)=2(2a+4)

OpenStudy (anonymous):

2a+4=56/2=28 2a=24 a=12

OpenStudy (sierra!):

ok thanks

OpenStudy (anonymous):

a is width so width is 12m